### v1

# Msmsfnet: a multi-stream and multi-scale fusion net for edge detection

Abstract
Edge detection is a long standing problem in computer vision. Recent deep learning based algorithms achieve state-of-the-art performance in publicly available datasets. Despite the efficiency of these algorithms, their performance, however, relies heavily on the pretrained weights of the backbone network on the ImageNet dataset. This limits heavily the design space of deep learning based edge detectors. Whenever we want to devise a new model, we have to train this new model on the ImageNet dataset first, and then fine tune the model using the edge detection datasets. The comparison would be unfair otherwise. However, it is usually not feasible for many researchers to train a model on the ImageNet dataset due to the limited computation resources. In this work, we study the performance that can be achieved by state-of-the-art deep learning based edge detectors in publicly available datasets when they are trained from scratch, and devise a new network architecture, the multi-stream and multi-scale fusion net (msmsfnet), for edge detection. We show in our experiments that by training all models from scratch to ensure the fairness of comparison, out model outperforms state-of-the-art deep learning based edge detectors in three publicly available datasets.
Index Terms:
Edge detection, deep learning, pretrained weights, ImageNet, msmsfnet.
publicationid: pubid:
I Introduction
Edge detection in both natural images [1, 2, 3, 4, 5, 6] and remote sensing images [7, 8, 9, 10, 11] has been studied for several decates due to the importance of edge features in image analysis and the use of edges in many applications such as line segment detection [12, 13, 14], image segmentation [5, 15], optical flow estimation [16] and object detection [17, 18]. Despite the huge amount of researches that are dedicated to the problem of edge detection, state-of-the-art edge detection algorithms mainly rely on deep learning. Convolutional Neural Networks (CNNs), due to its capability of learning multi-scale representations and hierarchical features, have been widely used in the field of computer vision. Recent deep learning based edge detectors fully exploit the merits of CNNs and constantly boot the performance of edge detection algorithms in publicly available datasets. In the early stage, the CNN-based edge detectors learned to detect edges from image patches [19, 20], until the proposal of the Holistically-nested edge detection  [21], which used a fully convolutional design and could thus be trained end-to-end and performed image-to-image prediction. Besides, the deep supervision technique used in HED enabled the network to learn and fuse multi-scale and multi-level features, which were important for extracting edges from various scales. Later on, various approaches were proposed following the fashion of HED, but with more complicated techniques to enhance the multi-scale feature learning and fusion cabability of the proposed methods. CEDN [22] trained a fully convolutional encoder-decoder network (the encoder was initialized using the VGG-16 network [23] and followed by a light-weight decoder network), in order to fully exploit the multi-scale representations learned by the pretrained VGG-16 net for object contour detection. Starting from the pretrained VGG net [23] and ResNet [24], COB [25] showed that the boundaries cound be detected more accurately by learning jointly the contour strength and contour orientations in multi-scale. RCF [26] proposed to combine the features extracted from all convolutional layers instead of using only the last layer of each stage as in HED, so as to fully exploit the multi-scale and multi-level information learned by all convolutional layers. By using more training data, the multi-scale testing technique and the improved network architecture, RCF achieved better performance than HED in the publicly available datasets for edge detection. CED [27] proposed a novel refinement network architecture, which increased the resolution of feature maps progressively with sub-pixel convolution, in order to improve the crispness of edges detected by HED. DeepBoundaries [28] extended HED with a multi-resolution architecture, and improved the edge detection accuracy of HED by using more training data and a well designed loss function. The performance of DeepBoundaries was further improved by combining the proposed multi-resolution architecture with the Normalized Cuts technique [29]. Similar to CED, the RefineContourNet (RCN) [30] was also designed following the idea of RefineNet [31], which used a multi-path refinement architecture, for more effective utilization of the high-level abstraction ability of the ResNet. BDCN [32] adopted a bi-directional cascade network structure, which supervised individual layers with edges at its specific scale, to extract edges at different scales. A Scale Enhancement Module (SEM) was also proposed in BDCN in order to enrich the multi-scale representations of each layer by using dilated convolutions. CATS [33] proposed to improve the localization accuracy of existing deep edge detectors with a novel tracing loss and a context-aware fusion block, tackling the problem of feature mixing and side mixing of existing methods.
Despite the efficiency of these end-to-end deep edge detectors, which constantly boosted the performance of edge detection, their performance, however, heavily relies on the pretrained weights of the backbone network on ImageNet dataset [34]. Apparently their performance will drop if the pretrained weights are not used, as much smaller amount of data is used for training. Apart from these methods, DexiNed [35] proposed a new network architecture for edge detection and created a new benchmark dataset (BIPEDv2).It showed that by training all models from scratch in the BIPEDv2 dataset, DexiNed outperformed all existing CNN-based edge detectors. Different from previous methods, DexiNed was not built on the basis of VGG net or the ResNet. Therefore, no pretrained weights on the ImageNet dataset is available for DexiNed, which could be an important reason why it does not provide quantitative comparisons on the previous boundary detection datasets.
As is known to all, the following four aspects are of crucial importance for the performance of a CNN model: the training data, network architecture, loss function and the training strategy (optimizers, hyperparameter settings). The use of pretrained weights on the ImageNet dataset in most existing CNN-based edge detectors makes it difficult to design new network architectures for edge detection, as we have to obtain an optimal pretrained weights for the new network by training it first using the ImageNet dataset, and then fine tune the model using the datasets for edge detection. However, it is quite expensive to train a model on ImageNet dataset, which contains more than one million images, and thus not feasible for many reserchers due to the limited computation resources. Besides, the pretrained weights of the VGG net or the ResNet might not help to improve the edge detection accuracy if the backbone network architecture is modified or our aim is to train the model for edge detection in a new kind of images, Synthetic Aperture Radar (SAR) images for example, because of the strong differences in the statistics between SAR and natural images [9]. Therefore, the use of pretrained weights in existing CNN-based edge detectors limits heavily the design space of CNN models for edge detection. To facilitate the design of new network architectures for edge detection in future work, we set new benchmarks to evaluate edge detection algorithms in publicly available datasets. We train state-of-the-art deep learning based edge detectors in three publicly available datasets for edge detection [36, 37, 35] and compute the optimal performance that can be achieved by them when training from scratch. We then propose a new network architecture for edge detection, the multi-stream and multi-scale fusion net (msmsfnet), and show that it outperforms these edge detectors when training from scratch in the three datasets.
The contributions of this work are two-fold:
  • we set new benchmarks to evaluate edge detection algorithms in publicly available datasets, which facilitates the design of new network architectures for edge detection;
  • we propose a new network architecture for edge detection, and it outperforms state-of-the-art edge detectors in three publicly available datasets when all models are trained from scratch.
The rest of the paper is organized as follows: In Section II, we provide the details about the proposed network architecture; in Section III, we demonstrate the efficiency of the proposed method by comparing it with state-of-the-art deep learning based edge detectors in three publicly available datasets for edge detection. Conclusions are finally given in Section IV.
II Msmsfnet: a multi-stream and multi-scale fusion net for edge detection
II-A Motivation
Developing deep convolutional neural networks for computer vision tasks has become popular ever since the proposal of Alexnet [38], which emphasizes the importance of depth on the performance of CNNs. Alexnet consists of five convolutional layers and three fully connected layers, among which kernels of large sizes ($11\times 1$ 11 and $5\times 5$) are used in the first two convolutional layers. Afterwards, the very deep convolutional networks (VGG net) was proposed in  [23], which utilized small convolutional filters ($3\times 3$) in all convolutional layers. As described in  [23], a sequence of $3\times 3$ convolutional filters has the same effective receptive field as convolutional filters of larger size, while it has fewer parameters to learn when the number of input and output channels is equal. The VGG net [23] showed that the accuracy of the network on the large scale image recognition tasks can be significantly boosted by pushing the depth of the network to 19 weight layers. Parallel to the work of VGG net, GoogLeNet [39] (also known as inceptionv1) achieved similarly high performance to the VGG net but with much more complicated network architecture. GoogLeNet pushed the depth of the network to 22 weight layers through the excessive use of dimension reduction with $1\times 1$ convolutions. Dimension reduction using $1\times 1$ convolution was also widely used in subsequent inception network architectures [40, 41, 42], the ResNet and its variants [43, 44], which reduces the number of parameters to be learned in each layer and thus makes it possible to increase the depth of the network significantly. Besides, the number of parameters in the network can also be reduced by factorizing large convolutional layers into small ones or by factorizing them into spatial asymmetric convolutions [41].
As demonstrated in previous works [38, 23, 39, 24, 40, 41, 42], the performance of CNN models can be boosted significantly by increasing the depth of the network, while a crucial factor to develop very deep CNNs is to reduce the number of parameters in each layer. In this work, we follow the strategy used in [41] and factorize usual convolutional filters into spatial asymmetric convolutions so as to reduce the number of parameters to be learned in each layer. Morespecifically, a $n\times n$ convolutional layer can be factorized into a $1\times n$ convolutional layer followed by a $n\times 1$ convolutional layer (or the reverse). This factorization reduces significantly the number of parameters to be learned while maintaining the size of the receptive field. For instance, factorizing a $3\times 3$ convolution into a $1\times 3$ convolution followed by a $3\times 1$ convolution contributes to a $33$ percent reduction in the number of parameters for the same number of input/output filters (assuming that the number of input and output filters is equal). Much more computational cost can be saved for the factorization of filters with larger kernel sizes. As what we will show in the following, we push the depth of our proposed model to 74 weight layers by utilizing excessively spatial asymmetric convolutions.
II-B Network architecture
For the task of edge detection, the multi-scale representation capability of models is of crucial importance. Existing CNN-based edge detectors constantly boost the performance of edge detection by enriching the multi-scale representations learned by the backbone networks in various ways. Although CNN models are naturally multi-scale/muti-level feature extractors because of the downsampled strided layers as well as the increasingly larger receptive field, we can still improve the multi-scale representation capability of CNNs by exploiting network architectures of higher quality.
Figure 1: The proposed msmsfblock. The expression in the parenthesis indicates the number of output filters of the layer.
To enrich the multi-scale representations, we devise a multi-stream and multi-scale fusion block (msmsfblock), and construct our msmsfnet by simply stacking our proposed msmsfblock. As shown in Figure 1 which displays the configurations of our msmsfblock, the input feature maps are fed into four parallel branches, extracting features at different scales due to the difference in the size of the receptive field. The extracted features are then fused by a two-step fusion stage, where the first step fuses features of closer scales, and the second step produces the final output of the msmsfblock. In our msmsfblock, we utilize excessively spatial asymmetric convolutions ($1\times n$ and $n\times 1$ convolutions), and only the final $3\times 1$ convolutional filter is followed by a ReLU function [38]. Batch normalization [40] is not used in our msmsfblock.
Our proposed msmsfnet is built by stacking the msmsfblock, with spatial max-pooling layers (kernel size $3\times 3$, stride 2) injected in-between. Similar to HED [21], we adopt the deep supervision technique [45]. We produce side outputs by adding side layers to the msmsfblocks before max-pooling layers and the final msmsfblock. In the side layers, a $3\times 3$ convolution is used to reduce the channel dimension of the feature maps to 1. If the feature maps have reduced resolution due to the downsampling operation done by the max-pooling layer, a deconvolutional layer which is fixed to perform bilinear interpolation is used to upsample the feature maps to the resolution of the input image. The final output of our msmsfnet are then obtained by fusing these side outputs with a $3\times 3$ convolutional layer. All side outputs and the fused output are supervised by the edge ground truth with a weighted cross entropy loss as defined in  [26]. Details about our proposed msmsfnet are displayed in figure 2.
Figure 2: Network architecture of the proposed msmsfnet. The number in the parenthesis indicates the number of output filters of the layer.
II-C Formulation
II-C1 Training phase
Given an image $u$ in the training dataset, notating $\mathbb{G}$ as its associated edge ground truth, where $u=\{u_{j},j=1,\cdots,|u|\}$ denotes the input image and $\mathbb{G}=\{\mathbb{G}_{j},j=1,\cdots,|\mathbb{G}|\}$ denotes the edge ground truth. Notating $\mathbb{W}$ the collection of parameters in all standard network layer (exculding those corresponding to side output layers and the final fusion layer), and notating $\mathbb{W}_{side}=(\mathbb{W}_{side}^{(1)},\mathbb{W}_{side}^{(2)},\mathbb{W}_{side}^{(3)})$ the collection of parameters in those side output layers, the objective function for the side output layers is defined as:
$\mathcal{L}_{side}(\mathbb{W},\mathbb{W}_{side})=\sum_{m=1}^{3}\beta_{m}\ell_{side}^{(m)}(\mathbb{W},\mathbb{W}_{side}^{(m)}),$ (1)
where $\ell_{side}$ denotes the loss corresponding to the side outputs.
In order to balance the loss between positive/negative classes, a class-balanced cross-entropy loss function as defined in  [26] is used in formula (1):
$\ell_{side}^{(m)}(\mathbb{W},\mathbb{W}_{side}^{(m)})$ $=$
$-\lambda\sum_{j\in\mathbb{G}_{+}}\log\mathbb{P}(\mathbb{G}_{j}=1|u:\mathbb{W},\mathbb{W}_{side}^{(m)})$
$-1.1\times(1-\lambda)\sum_{j\in\mathbb{G}_{-}}\log\mathbb{P}(\mathbb{G}_{j}|u:\mathbb{W},\mathbb{W}_{side}^{(m)}),$ (2)
where $\lambda=\frac{|\mathbb{G}_{-}|}{|\mathbb{G}|}$,and $1-\lambda=\frac{|\mathbb{G}_{+}|}{|\mathbb{G}|}$, $|\mathbb{G}_{-}|$ and $|\mathbb{G}_{+}|$ represent the non-edge and edge label sets. $\mathbb{P}(\mathbb{G}_{j}=1|u:\mathbb{W},\mathbb{W}_{side}^{m})=\sigma(a_{j}^{(m)})\in[0,1]$ with $\sigma(\cdot)$ representing the sigmoid function and $a_{j}^{(m)}$ representing the activation value at pixel $j$. The edge probability map produced by each side layer is thus defined as:
$\hat{\mathbb{G}}_{side}^{(m)}=\sigma(\hat{A}_{side}^{m}),$ (3)
where $\hat{A}_{side}^{(m)}\equiv\{a_{j}^{(m)},j=1,2,\cdots,|\mathbb{G}|$ represent the activations of the output of the side layer.
The final fused output of the model is obtained by fusing these side outputs with a $3\times 3$ convolutional layer. The same loss function as the side layers is used to compute the loss between the edge ground truth and the final fused output.
Denoting $\mathbb{W}_{fuse}$ the collection of parameters corresponding to the fusion layer and $\mathcal{L}_{fuse}(\cdot)$ the loss function corresponding to the fused output, the objective function that needs to be minimized during training is thus defined as:
$(\mathbb{W},\mathbb{W}_{side},\mathbb{W}_{fuse})^{\star}$ $=$
$\operatorname*{argmin}(\mathcal{L}_{side}(\mathbb{W},\mathbb{W}_{side})+\mathcal{L}_{fuse}(\mathbb{W},\mathbb{W}_{side},\mathbb{W}_{fuse})).$ (4)
II-C2 Testing phase
Given a testing image $u$, four edge probability maps (three side outputs and one fused output) are obtained from our msmsfnet:
$(\hat{\mathbb{G}}_{fuse},\hat{\mathbb{G}}_{side}^{(1)},\cdots,\hat{\mathbb{G}}_{side}^{(3)})$ $=$
$\mathrm{CNN}(u,(\mathbb{W},\mathbb{W}_{side},\mathbb{W}_{fuse})^{\star}).$ (5)
The fused output is chosen as the final output of our network.
III Experiments
In this section, we demonstrate the efficiency of our proposed msmsfnet by comparing it with several representative CNN-based edge detectors (HED [21], RCF [26], BDCN [32], CATS [33], DexiNed [35]) in three publicly available datasets for edge detection (the BIPEDv2 dataset [35], the BSDS500 dataset [36] and the NYUDv2 dataset) [37]. We show in our experiments that our method achieves better performance than state-of-the-art CNN-based edge detectors in all these datasets when all models are trained from scratch.
III-A Implementation details
The proposed msmsfnet is implemented using the PyTorch library, and trained from scratch using the Adam optimizer [46]. The weights of all convolutional layers are randomly initialized using the Xavier uniform distribution and the biases are all initialized to 0. Weight decay is set to $10^{-12}$. The mini-batch size for training is set to 6. The learning rate is initialized to $10^{-4}$, and decreased by 10 after a fixed number of epochs (10 for the BIPEDv2 dataset, 20 for the BSDS500 dataset, and 20 for the NYUDv2 dataset). The maximum numner of epochs to train is set to 15 for the BIPEDv2 dataset, and 25 for the BSDS500 and NYUDv2 dataset. The optimal performance obtained by each method during training is reported in the following.
III-B Evaluation criteria
The outputs of all CNN-based edge detectors are edge probability maps instead of binary edge maps. Non-maxima suppression, thresholding and edge thining are still required to obtain binary edge maps. Following previous works [21, 26, 32], we adopt the non-maxima suppression step of  [6] to process the outputs of our model. Three criteria are used to compare the performance of different methods: the Optimal Dataset Scale (ODS) F1-score, where a global optimal threshold is chosen to binarize the edge probability maps; the Optimal Image Scale (OIS) F1-score, where an optimal threshold is chosen for each image to binarize the edge probability maps ; and the Average Precision (AP). Higher values of these criteria indicate better performance obtained by a method. Following previous works, precision-recall curves are also ploted to compare the performance of different methods. In the following, we will compare the performance of our msmsfnet with state-of-the-art deep learning based edge detectors using these criteria in three publicly available datasets.
III-C BIPEDv2 dataset
We first compare the performance of our proposed msmsfnet with HED [21], RCF [26], BDCN [32], CATS [33] and DexiNed [35] in the second version of the Barcelona Images for Perceptual Edge Detection (BIPEDv2) dataset [35], as the performance reported by existing algorithms on it does not involve the use of pretrained weights on the ImageNet dataset.
TABLE I: Quantitative comparison of different algorithms on the 50 testing images of the BIPEDv2 dataset.
Methods ODS F1-score OIS F1-score AP
HED 0.884 0.890 0.919
RCF 0.882 0.889 0.911
BDCN 0.893 0.900 0.922
CATS 0.894 0.899 0.927
DexiNed 0.894 0.900 0.921
msmsfnet 0.897 0.901 0.936
The BIPEDv2 dataset contains 200 images for training and 50 images for testing. All images are outdoor images and of size $1280\times 720$ ppixels. We use the data augmentation strategy described in  [35] to augment the training dataset, and train all all models from scratch using the augmented training dataset. Quantitative comparisons of different methods on the 50 testing images of the BIPEDv2 dataset are displayed in table I.
From table I we can see that the proposed msmsfnet obtains the highest scores in all three criteria. Minor improvements over existing methods (0.3 percent for ODS F1-score and 0.1 percent for OIS F1-score) are made by our msmsfnet on the test set of BIPEDv2 dataset. The improvents made by our method are relatively larger for the Average Precision (at least 0.9 percent). The precision-recall curves computed by different methods in the test set of BIPEDv2 dataset are displayed in figure 3.
As shown in table I and figure 3, the proposed msmsfnet achieves slightly better performance than existing methods in the BIPEDv2 dataset. This is probably because the edge anotations in the dataset are ’low level edges’, while the multi-representation capability of a model plays a more important role in higher level tasks such as boundary/contour detection. As what we will show in follwing, much larger improvements are obtained by the proposed msmsfnet in the boundary/contour detection datasets. Details about the differences among edges, boundaries, and contours can be found in  [35] (following previous works [21, 26, 25, 32, 33], we do not strictly distinguish between edge and boundary/contour detection and call all these tasks as edge detection).
Figure 3: The precision-recall curves computed by different methods in the test set of the BIPEDv2 dataset.
The edge maps computed by different methods in an image from the test set of BIPEDv2 dataset are displayed in figure 4. From figure 4 we can see that the edge maps computed by HED, RCF and DexiNed are noisy. Many false edges are detected by them. Compared to BDCN and CATS, fewer false edges are detected by our proposed method. Compare to the edge maps computed by other methods, the edge map computed by our method is cleaner and closer to the edge ground truth.
Refer to caption Refer to caption
(a) image (b) ground truth
Refer to caption Refer to caption
(c) HED (d) RCF
Refer to caption Refer to caption
(e) BDCN (f) CATS
Refer to caption Refer to caption
(g) DexiNed (h) msmsfnet
Figure 4: Edge detection results computed by different edge detectors in a test image of BIPEDv2 dataset.
III-D BSDS500 dataset
[…]
TABLE II: Quantitative comparisons of different algorithms on the test set of BSDS500 dataset.
Methods ODS F1-score OIS F1-score AP
HED 0.798 0.815 0.825
RCF 0.798 0.815 0.815
BDCN 0.803 0.821 0.837
CATS 0.801 0.820 0.841
DexiNed 0.799 0.818 0.810
msmsfnet 0.816 0.835 0.859
Table II shows the quantitatie comparisons of different methods in the test set of BSDS500 dataset. It is clearly shown that the proposed msmsfnet achieves much higher scores than state-of-the-art edge detectors under the same experimental settings. The ODS F1-score achieved by our methods is at lease 1.3 higher percent than existing methods. In particular, the ODS F1-score computed by our method is 1.3 percent higher than humans, which shows that the perception ability of our model is stronger than humans even without the need to pretrain the model on the ImageNet dataset. The Average Precision computed by our method is at lease 1.8 percent higher than existing methods. The performance gains obtained by our method in the BSDS500 dataset indicate the strong multi-scale representation capability of our model. Precision-recall curves computed by differenet methods are displayed in figure 5. The edge maps computed by different methods in an image from the test set of BSDS500 dataset are shown in figure 6. It is clearly shown that the edge map computed by our method is closer to the edge ground truth. More true edges are detected with fewer false detections. In contrast, the edges detected by BDCN, CATS and DexiNed tend to be fragmented, while many false edges are detected by HED.
Figure 5: The precision-recall curves computed by different methods in the test set of BSDS500 dataset.
Refer to caption Refer to caption
(a) image (b) ground truth
Refer to caption Refer to caption
(c) HED (d) RCF
Refer to caption Refer to caption
(e) BDCN (f) CATS
Refer to caption Refer to caption
(g) DexiNed (h) msmsfnet
Figure 6: Edge detection results computed by different methods in a test image of BSDS500 dataset.
III-E NYUDv2 dataset
The NYU-Depth V2 (NYUDv2) dataset [37] contains 1449 pairs of RGB and depth images, which are well aligned, densely labeled and made into the same size. These images have been splitted into three parts: 381 images for training, 414 images for validation and 654 images for testing, The images in the dataset are video sequences of various indoor scenes and recorded by both RGB and Depth camera from Microsoft Kinect. Following previous works [26, 32], both training and validation images are augmented and used to train the models. We adopt the same data augmentation strategy as  [26]. Each model has been trained twice: one trained using RGB images and one trained using HHA features. We compare the performance of different methods using three types of edge predictions produced by each model: one produced by the model trained on RGB images, one produced by the model trained on HHA features, and one produced by averaging the two previous predictions. During evaluation, the maximum tolerance allowed to math the edge predictions with the edge ground truth correctly is increased from 0.0075 to 0.011 as done previous works [21, 26, 32]. Quantitative comparisons of different methods in the test set of NYUDv2 dataset are displayed in table III.
TABLE III: Quantitative comparisons of different algorithms on the test set of NYUDv2 dataset.
Methods ODS F1-score OIS F1-score AP
$\text{HED}_{\text{RGB}}$ 0.704 0.720 0.714
$\text{RCF}_{\text{RGB}}$ 0.715 0.731 0.717
$\text{BDCN}_{\text{RGB}}$ 0.709 0.728 0.692
$\text{CATS}_{\text{RGB}}$ 0.713 0.730 0.721
$\text{DexiNed}_{\text{RGB}}$ 0.721 0.734 0.650
$\text{msmsfnet}_{\text{RGB}}$ 0.732 0.747 0.744
$\text{HED}_{\text{HHA}}$ 0.674 0.688 0.680
$\text{RCF}_{\text{HHA}}$ 0.685 0.701 0.680
$\text{BDCN}_{\text{HHA}}$ 0.677 0.694 0.692
$\text{CATS}_{\text{HHA}}$ 0.678 0.695 0.679
$\text{DexiNed}_{\text{HHA}}$ 0.669 0.686 0.612
$\text{msmsfnet}_{\text{HHA}}$ 0.704 0.721 0.727
$\text{HED}_{\text{RGB+HHA}}$ 0.716 0.739 0.758
$\text{RCF}_{\text{RGB+HHA}}$ 0.730 0.753 0.764
$\text{BDCN}_{\text{RGB+HHA}}$ 0.728 0.751 0.774
$\text{CATS}_{\text{RGB+HHA}}$ 0.726 0.748 0.773
$\text{DexiNed}_{\text{RGB+HHA}}$ 0.743 0.757 0.735
$\text{msmsfnet}_{\text{RGB+HHA}}$ 0.746 0.767 0.789
Refer to caption Refer to caption Refer to caption
(a) RGB (b) HHA (c) RGB$+$HHA
Figure 7: The precision-recall curves computed by different methods in the NYUDv2 dataset.
As shown in table III, our model achieves the highest scores in all experimental settings. When all models are trained using only RGB images, the ODS F1-score computed by our method is at least 1.1 percent higher than those computed by existing methods. The Average Precision computed by our method is at least 2.3 percent higher than those computed by existing methods. When only HHA features are used to train the model, the ODS F1-score computed by our method is at least 1.9 percent higher. The improvements on the Average Precision is much larger, at least 3.5 percent higher than state-of-the-art methods. For the edge predictions computed by combining the information from both RGB images and HHA features, the improvements on the ODS F1-score is relatively minor, only 0.3 percent higher than the best result computed by DexiNed. However, the Average Precision computed by our method is 5.4 percent higher than that of DexiNed. The ODS F1-score computed by our method is at least 1.6 percent higher than the rest methods. Similar improvements have been done on the Average Precision, at least 1.5 percent higher than the rest methods. The precision-recall curves computed by different methods in all experimental settings are displayed in figure 7. The edge maps computed by different methods in an image from the test set of the NYUDv2 dataset can be found in figure 8, figure 9 and figure 10. Figure 8 displays the edge detection results computed by different methods using only RGB images. Figure 9 displays the results computed different methods using only HHA features, while figure 10 displays the results computed using the information from both RGB images and HHA features. It can be seen from these figures that the edges detected by our method are more complete and fewer false edges are detected.
Refer to caption Refer to caption
(a) RGB image (b) ground truth
Refer to caption Refer to caption
(c) HED (d) RCF
Refer to caption Refer to caption
(e) BDCN (f) CATS
Refer to caption Refer to caption
(g) DexiNed (h) msmsfnet
Figure 8: Edge detection results computed by different methods in a test image (RGB) of the NYUDv2 dataset.
Refer to caption Refer to caption
(a) HHA feature (b) ground truth
Refer to caption Refer to caption
(c) HED (d) RCF
Refer to caption Refer to caption
(e) BDCN (f) CATS
Refer to caption Refer to caption
(g) DexiNed (h) msmsfnet
Figure 9: Edge detection results computed by different methods in a test image (HHA) of the NYUDv2 dataset.
Refer to caption Refer to caption
(a) RGB image (b) HHA feature
Refer to caption Refer to caption
(c) HED (d) RCF
Refer to caption Refer to caption
(e) BDCN (f) CATS
Refer to caption Refer to caption
(g) DexiNed (h) msmsfnet
Figure 10: Edge detection results computed by different methods using both RGB and HHA features of a test image in the NYUDv2 dataset.
IV Conclusions
In this paper we focus on the task of edge detection in natural images. Unlike previous works that use pretrained ImageNet models, in this work all models are trained from scratch and evaluated under the same experimental settings. By doing so, we ensure the fairness to compare the detection capability of different models. In the meantime, the experimental settings in this work facilitate the design of new network architectures for edge detection, as pretraining a model on the ImageNet dataset is not pratical for many researchers due to the limited computation resources. Besides, we propose a new model for edge detection, which adopts a multi-branch architecture to enhance its multi-scale representation capability and utilizes spatial asymmetric convolutions so as to increase its accuracy by developing a very deep model. The proposed msmsfnet achieves superior performance than state-of-the-art methods in three publicly available datasets under the same experimental settings, which shows the efficiency of the proposed method.
Source code
The source code of the method will be released upon the acceptance of the paper .
Fundings
This work is supported in part by the National Natural Science Foundation of China (Grant No. 42374018), in part by the Shenzhen Scientific Research and the Development Funding Program (Grant No. RCYX20210706092140076), in part by the Swift Fund Fintech Funding and the Open Research Fund from Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ, Grant No. GML-KF-22-15), in part by the Shenzhen Science and Technology Program (Grant No. JCYJ20220531101409021), and in part by the National Natural Science Foundation of China (Grant No. 62371348).